OpenStudy (anonymous):

solve and check 2x^2+6x=8

OpenStudy (nathan917):

2x^2+6x=8 Simple and best practice solution for 2x^2+6x=8 equation. Check how easy it is, and learn it for the future. Our solution is simple, and easy to understand, so don`t hesitate to use it as a solution of your homework. If it's not what You are looking for type in the equation solver your own equation and let us solve it. Equation: Solution for 2x^2+6x=8 equation: Simplifying 2x2 + 6x = 8 Reorder the terms: 6x + 2x2 = 8 Solving 6x + 2x2 = 8 Solving for variable 'x'. Reorder the terms: -8 + 6x + 2x2 = 8 + -8 Combine like terms: 8 + -8 = 0 -8 + 6x + 2x2 = 0 Factor out the Greatest Common Factor (GCF), '2'. 2(-4 + 3x + x2) = 0 Factor a trinomial. 2((-4 + -1x)(1 + -1x)) = 0 Ignore the factor 2. Subproblem 1 Set the factor '(-4 + -1x)' equal to zero and attempt to solve: Simplifying -4 + -1x = 0 Solving -4 + -1x = 0 Move all terms containing x to the left, all other terms to the right. Add '4' to each side of the equation. -4 + 4 + -1x = 0 + 4 Combine like terms: -4 + 4 = 0 0 + -1x = 0 + 4 -1x = 0 + 4 Combine like terms: 0 + 4 = 4 -1x = 4 Divide each side by '-1'. x = -4 Simplifying x = -4 Subproblem 2 Set the factor '(1 + -1x)' equal to zero and attempt to solve: Simplifying 1 + -1x = 0 Solving 1 + -1x = 0 Move all terms containing x to the left, all other terms to the right. Add '-1' to each side of the equation. 1 + -1 + -1x = 0 + -1 Combine like terms: 1 + -1 = 0 0 + -1x = 0 + -1 -1x = 0 + -1 Combine like terms: 0 + -1 = -1 -1x = -1 Divide each side by '-1'. x = 1 Simplifying x = 1 Solution x = {-4, 1}

OpenStudy (anonymous):

I know that I have to -8 to set the problem to equal 0

OpenStudy (nathan917):

That a bit much? if so i can cut it down

OpenStudy (anonymous):

no its perfect thank you

OpenStudy (nathan917):

Np.